# Exploring the Effects of Load Altering Attacks on Load Frequency Control through Python and RTDS

### v1

@@ II-B LAA Modeling
SLAAs can be incorporated into the LFC model as they only affect $\Delta P_{L}$. This additively affects $\Delta f$ by $\frac{-\Delta P_{L}}{2H}$, influencing all the other variables dependent on such deviation. Alternatively, SLAAs can be modeled as a set of differential equations [18]. In this case, as shown in (16), the system’s internal state is modeled as a concatenation of vectors $\delta$, $\theta$, and $\omega$: voltage phase angles of generator buses, voltage phase angles of load buses, and rotor frequency deviation of generator buses. The vector $P^{LS}$ defines the system input as a secure portion of the controllable frequency-sensitive load at each bus. In contrast, vector $\epsilon^{L}$ represents the vulnerable and compromised portion of the load used to perform SLAA.
@@
The (17), and (18) show the system, input, and mass matrices $A$, $B$, and $E$. The diagonal matrix $M$ shows the generators’ inertia. $D^{G}$ is the diagonal matrix of the generator damping coefficients. The diagonal matrices $K^{P}$ and $K^{I}$ show the proportional and integral coefficients for primary and secondary control. The transmission lines susceptance matrix $H_{bus}$ (18) depicts connections between generator-to-generator ($H^{GG}$), generator-to-load bus ($H^{GL}$), load bus-to-generator ($H^{LG}$), and load bus-to-load bus ($H^{LL}$). If two buses are unconnected, the respective matrix element equals zero. Lastly, $I$ is the identity matrix of the appropriate dimensions.
$A=\begin{bmatrix}0&0&I\\ H^{LG}&H^{LL}&0\\ K^{I}+H^{GG}&H^{GL}&K^{P}+D^{G}\end{bmatrix};\quad B=\begin{bmatrix}0\\ I\\ 0\end{bmatrix}$ (17)
$E=\begin{bmatrix}I&0&\kern-4.5pt0\\ 0&0&\kern-4.5pt0\\ 0&0&\kern-4.5pt-M\end{bmatrix};\quad H_{bus}=\begin{bmatrix}H^{GG}H^{GL}\\ H^{LG}H^{LL}\end{bmatrix}$ (18)
While the previous model effectively represents SLAA, this attack does not alter the system stability, a key distinction of DLAA [8]. To model DLAA, we modify (16), incorporating the attack into the system matrix to affect stability. As shown in (19), the DLAA proportional coefficient vector $K^{LG}>\nolinebreak 0$ indicates the vulnerable load portion that can modify the system state during an attack.
$E\begin{bmatrix}\dot{\delta}\\ \dot{\theta}\\ \dot{\omega}\end{bmatrix}=A\begin{bmatrix}\delta\\ \theta\\ \omega\end{bmatrix}+B\left(\begin{bmatrix}0\\ 0\\ -K^{LG}\end{bmatrix}^{T}\begin{bmatrix}\delta\\ \theta\\ \omega\end{bmatrix}+P^{LS}\right)$ (19)
The system matrix also changes as in (20). By substituting $\theta$ (21) and transforming state-space representation into a non-descriptor form, we obtain the new system matrix $A^{*}$ (22) of the system under attack. The new form of the system state-space equations is shown in (23). This model integrates the DLAA into the system matrix, meaning that changes to $K^{LG}$ can affect system stability by shifting the eigenvalues of $A^{*}$.
$A^{*}=A+B\begin{bmatrix}0&0&\kern-3.59995pt-K^{LG}\end{bmatrix}$ (20)
$\kern 3.59995pt\theta\kern 3.96002pt=(H^{LL})^{-1}K^{LG}\omega-(H^{LL})^{-1}H^{LG}\delta-(H^{LL})^{-1}P^{LS}$ (21)
@@
$\begin{bmatrix}\dot{\delta}\\ \dot{\omega}\end{bmatrix}=A^{*}\begin{bmatrix}\delta\\ \omega\end{bmatrix}+\begin{bmatrix}0\\ -(H^{LL})^{-1}\end{bmatrix}P^{LS}$ (23)

### v2

@@ II-B LAA Modeling
SLAAs can be incorporated into the LFC model as they only affect $\Delta P_{L}$. This additively affects $\Delta f$ by $\frac{-\Delta P_{L}}{2H}$, influencing all the other variables dependent on such deviation. Alternatively, SLAAs can be modeled as a set of differential equations [18]. In this case, as shown in (16), the system’s internal state is modeled as a concatenation of vectors $\delta$, $\theta$, and $\omega$: voltage phase angles of generator buses, voltage phase angles of load buses, and rotor frequency deviation of generator buses. The vector $P^{LS}$ defines the system input as a secure portion of the load at each bus. In contrast, vector $\epsilon^{L}$ represents the vulnerable portion of the load used to perform SLAA.
@@
The (17), and (18) show the system, input, and mass matrices $A$, $B$, and $E$. We define $A^{\kern-0.70004pt1}\!\!:=\!diag(\!A\!*\!\mathbf{1}\!)$ where $\mathbf{1}$ is the column vector of ones. The diagonal matrices $M$, $D^{G}$, $K^{P}$, and $K^{I}$ show the generator inertias, damping coefficients, and proportional and integral coefficients for primary and secondary control accordingly. The admittance matrix $H_{bus}$ (18) depicts connections between generator-to-generator ($H^{GG}$), generator-to-load bus ($H^{GL}$), load bus-to-generator ($H^{LG}$), and load bus-to-load bus ($H^{LL}$). If two buses are unconnected, the respective matrix element equals zero. Lastly, $I$ is the identity matrix of the appropriate dimensions.
$A=\begin{bmatrix}0&0&I\\ -\!H^{\!LG}&H^{\!LG^{1}}\kern-5.40005pt+\!\!H^{\!LL^{1}}\kern-6.29997pt-\!\!H^{\!LL}&0\\ K^{\!I}\kern-3.59995pt+\!\!H^{\!GG^{1}}\kern-6.29997pt-\!\!H^{\!GG}\kern-3.59995pt+\!\!H^{\!GL^{1}}&-\!H^{\!GL}&K^{\!P}\kern-4.5pt+\!\!D^{\kern-0.44104ptG}\end{bmatrix}$ (17)
$E=\begin{bmatrix}I&0&\kern-4.5pt0\\ 0&0&\kern-4.5pt0\\ 0&0&\kern-4.5pt-M\end{bmatrix};\quad B=\begin{bmatrix}0\\ I\\ 0\end{bmatrix};\quad H_{bus}=\begin{bmatrix}H^{GG}H^{GL}\\ H^{LG}H^{LL}\end{bmatrix}$ (18)
While the previous model effectively represents SLAA, this attack does not alter the system stability, a key distinction of DLAA [8]. To model DLAA, first we modify (16) by substituting $\theta$ (19) to then obtain the non-descriptor form. Next, we incorporate the attack into the system matrix to affect stability. As shown in (20), the DLAA proportional coefficient vector $K^{LG}>\nolinebreak 0$ indicates manipulation of the vulnerable load portion that can modify the system state during an attack.
$\theta=H^{inv}(H^{LG}\delta\!-\!P^{L});\quad H^{inv}\kern-1.79997pt=(H^{LG^{1}}\kern-3.59995pt+\!H^{LL^{1}}\kern-4.5pt-\!H^{LL})^{-1}$ (19)
$\begin{bmatrix}\dot{\delta}\\ \dot{\omega}\end{bmatrix}=A^{\prime}\begin{bmatrix}\delta\\ \omega\end{bmatrix}+B^{\prime}\left(\begin{bmatrix}0&\kern-3.59995pt-K^{LG}\end{bmatrix}\begin{bmatrix}\delta\\ \omega\end{bmatrix}+P^{LS}\right)$ (20)
The system matrix also changes as in (21). We obtain the new system matrix $A^{*}$ (22) of the system under attack. The new form of the system state-space equations is shown in (23). This model integrates the DLAA into the system matrix, meaning that changes to $K^{LG}$ can affect system stability by shifting the eigenvalues of $A^{*}$.
$A^{*}=A^{\prime}+B^{\prime}\begin{bmatrix}0&\kern-3.59995pt-K^{LG}\end{bmatrix}$ (21)
@@
$\begin{bmatrix}\dot{\delta}\\ \dot{\omega}\end{bmatrix}=A^{*}\begin{bmatrix}\delta\\ \omega\end{bmatrix}+\begin{bmatrix}0\\ M^{-1}H^{GL}H^{inv}\end{bmatrix}P^{LS}$ (23)